OpenStudy (anonymous):

Expand each Logarithm...? log square root x square root 2 divided by y^2 log base 3 [(xy)1/3]^3 log (xyz)^1/2 divided by z^2

OpenStudy (anonymous):

So, in general if you have something like: \[\ln \left[ \left( \frac{a * b}{c} \right)^{d} \right] = d \ln \left[ \left( \frac{a * b}{c} \right)\right]= d \left[ \ln (a * b) - \ln(c)\right] = d \left[ \ln(a)+\ln(b)-\ln(c)\right]\] Using these properties: \[\ln(a*b)=\ln(a)+\ln(b); \ln(a/b)= \ln (a) - \ln(b); \ln(a^b)=b \ln(a)\] Now try to apply this to your problems. I can help you more if needed.

OpenStudy (anonymous):

Could you solve one problem?

OpenStudy (anonymous):

\[\ln(\sqrt{2}\sqrt{x}/y^2)=\ln(\sqrt{2x})-\ln(y^2)=\frac{1}{2}\ln(2x)-2\ln(y)=\frac{1}{2}(\ln(2)+\ln(x))-2\ln(y)\]

OpenStudy (anonymous):

-2lny** at the end.

OpenStudy (anonymous):

why is it so complicated?

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

No problem.